make a sketch that matches the description: B is between A and D. C is between B and D. AB= 7 cm, AC= 15 cm, and AD= 32 cm. Find BC, BD, and CD.
1
Expert's answer
2012-09-11T09:27:12-0400
A_______B________C_________________D
& BC=AC-AB=15-7=8 cm
BD=AD-AB=32-7=25 cm
CD=AD-AC=32-15=17 cm